OpenStudy (anonymous):

Let X be Unif(0,theta) consider estimators of the form theta_hat = aX where a>0 is an arbitrary constant, For what a is theta_hat unbiased?

OpenStudy (zarkon):

do you know what \(\mathbb{E}(X)\) is?

OpenStudy (anonymous):

Zarkon no

OpenStudy (zarkon):

\[\int\limits_{0}^{\theta}x\frac{1}{\theta}dx\]

OpenStudy (anonymous):

Zarkon would you have time explaining me couple of stuff in statistics?

OpenStudy (anonymous):

where did the 1/theta come from?

OpenStudy (zarkon):

that is the pdf of a uniform(0,theta)